OpenStudy (anonymous):

Solve 3abs(lnx)-12=0

OpenStudy (anonymous):

\[3\left| \ln x \right|-12=0\]

OpenStudy (anonymous):

@jim_thompson5910 Could you help me again?

jimthompson5910 (jim_thompson5910):

\[\Large 3|\ln(x)| - 12=0\] \[\Large 3|\ln(x)|=12\] \[\Large |\ln(x)|=\frac{12}{3}\] \[\Large |\ln(x)|=4\] \[\Large \ln(x)=4 \ \text{ or } \ \ln(x)=-4\] I'll let you finish

jimthompson5910 (jim_thompson5910):

keep in mind that \[\Large \ln(x) = \log_{e}(x)\] 'e' is like pi and it is approximately 2.718 (the decimal expansion goes on forever with no known pattern)

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

np